# Dynamics of Cosmic Superstrings and the Overshoot Problem

### v1

@@ I Introduction
Early universe cosmology is a very promising arena to probe UV complete theories like string theory. This involves the study of trans-Planckian field excursions during epochs of large field inflation and kination, or the gravitational decay of long-lived moduli (see Cicoli et al. (2024) for a recent review of different aspects of string cosmology).
@@
In this paper we focus on two features that typically arise at the end of string inflation (see for example brane-antibrane inflation Burgess et al. (2001); Dvali et al. (2001); Kachru et al. (2003a); Baumann et al. (2008, 2007); Cicoli et al. (2025) or volume modulus inflation Conlon et al. (2008); Cicoli et al. (2015)): a volume modulus that rolls over several Planck units towards its late time minimum, and a population of cosmic superstrings with a modulus-dependent tension that varies with time.
@@
A notorious issue of string models of the early universe is the overshoot problem Brustein and Steinhardt (1993), that is the tendency of the volume mode to run over the barrier that separates its minimum from the decompactification limit, if the modulus is released with a very high energy after the end of inflation. A known solution relies on the presence of a background fluid, like radiation or matter, which redshifts slower than the modulus kinetic energy and yields additional Hubble friction to the modulus dynamics Kaloper and Olive (1993); Barreiro et al. (1998); Brustein et al. (2004); Kaloper et al. (2005); Battefeld and Watson (2006); Conlon et al. (2008); Acharya et al. (2008). However, especially in models where the volume mode is rolling before reheating, motivating the microscopic origin of enough radiation to prevent overshooting can be non-trivial Conlon and Revello (2022); Mosny et al. (2025).
@@
Recently, the powerful techniques of dynamical systems have been exploited to explore in detail the cosmological evolution of systems involving rolling scalars, cosmic superstrings and additional background fluids. Previous studies discovered that, during kination, the physical size of loops of fundamental strings (hereafter referred to as F-strings) Conlon et al. (2024) and effective strings arising from D3- and NS5-branes wrapped around fibration cycles Brunelli et al. (2025) can grow faster than the scale factor and form a cosmic superstring network. Interestingly, this growth can occur for F-strings also for scaling fixed points in the presence of a background fluid with an appropriate equation of state parameter Brunelli et al. (2025). Moreover, ref. Sánchez González et al. (2025) took into account the energy density in cosmic string loops and revealed, among other things, the existence of a new attractor fixed point (a loop tracker) where the energy density of the universe is dominated by the cosmic superstrings.
@@
In this paper we extend the analysis of these cosmological systems considering the whole potential of the rolling volume modulus of type IIB LVS scenarios Balasubramanian et al. (2005); Cicoli et al. (2008), which includes both an early-time runaway and a late-time minimum with a barrier against decompactification. We find that, without enougth radiation, a model with F-strings would face an overshoot problem. On the other hand, as soon as an initial population of effective strings arising from NS5-branes wrapped around 4-cycles is present, the overshoot problem is avoided, even in the absence of radiation. Moreover, the cosmic superstrings constitute around $97\%$ of the total energy density in the loop tracker fixed point that precedes the final evolution of the field towards the minimum of its potential where the energy density of cosmic superstrings can be as high as about $50\%$. Such large values of energy densities imply the potential generation of a detectable gravitational wave (GW) signal at high frequencies from the decay of cosmic superstrings in the early universe.
II Strings with time-dependent tension
@@
We start by giving a brief review of the derivation of the equations of motion for a string loop with time-dependent tension following Emond et al. (2022); Conlon et al. (2024); Revello and Villa (2025). The Nambu-Goto (NG) action for a string with a tension $\mu$ dependent on the spacetime coordinates $x^{\mu}$ is given by:
$S_{\rm NG}=-\int d^{2}\sigma\,\mu(x)\,\sqrt{-\gamma}\,.$ (1)
@@
$\gamma_{ab}=\frac{\partial x^{\mu}}{\partial\sigma^{a}}\frac{\partial x^{\nu}}{\partial\sigma^{b}}\,g_{\mu\nu}\,.$ (2)
Varying (1) with respect to $x^{\mu}$ yields the following equations of motion Emond et al. (2022):
$\frac{1}{\mu\sqrt{-\gamma}}\,\partial_{a}(\mu\sqrt{-\gamma}\gamma^{ab}x^{\mu}_{,b})+\Gamma^{\mu}_{\nu\rho}\gamma^{ab}x^{\nu}_{,a}x^{\rho}_{,b}-\frac{\partial^{\mu}\mu}{\mu}=0\,.$ (3)
@@
$\frac{\dot{\epsilon}}{\epsilon}$ $=H-a^{2}\dot{R}^{2}\left(2H+\frac{\dot{\mu}}{\mu}\right),$ (5)
$\ddot{R}+H\dot{R}\,+$ $\frac{R}{\epsilon^{2}}+\left(2H+\frac{\dot{\mu}}{\mu}\right)(1-a^{2}\dot{R}^{2})\dot{R}=0\,$, (6)
where $\epsilon\equiv a|R|/\sqrt{1-a^{2}\dot{R}^{2}}$ parametrizes the physical radius of the string at the oscillation turning points and is related to its (linear) energy density Kibble (1976).
@@
$\mu(t)=\mu_{0}\,e^{-\sqrt{6}\beta\,\frac{\phi(t)}{M_{p}}}\ .$ (7)
We will first employ the techniques of Sánchez González et al. (2025) to study the cosmological evolution of a system including a population of cosmic superstrings and a modulus rolling towards the minimum of its potential, with and without radiation in generality. Then we will study what happens when the field stabilises at its minimum and the tension becomes oscillating, analysing the evolution of a single isolated string and whether it has any peculiar dynamics.
III Evolution towards the minimum
All previous works on the evolution of a dilute gas of cosmic string loops assumed an exponential potential Conlon et al. (2024); Brunelli et al. (2025); Sánchez González et al. (2025). However this approximation holds only in a portion of a phenomenologically realistic potential. In particular, when the rolling field is the volume modulus of string compactifications, it must settle into a minimum at the end of its excursion to avoid a decompactification limit. The precise structure of the potential thus becomes relevant for assessing how the field reaches its minimum and whether it settles there without overshooting.
The potential we will focus on is the one of type IIB LVS models Balasubramanian et al. (2005); Cicoli et al. (2008). This class of models features (at least) two moduli fields: one controlling the overall internal volume, and the other parametrising the size of a small blow-up cycle. At leading order, the LVS potential receives three types of contributions: $\mathcal{O}(\alpha^{\prime 3})$ corrections to the Kähler potential Becker et al. (2002), non-perturbative effects in the superpotential Blumenhagen et al. (2009), and a model-dependent uplift term Kachru et al. (2003b); Cicoli et al. (2016, 2012). After integrating out the blow-up mode, the potential depends only on the volume mode $\mathcal{V}$. The canonical normalisation of the volume is:
$\frac{\Phi}{M_{p}}=\sqrt{\frac{2}{3}}\ln\mathcal{V}\,.$ (8)
@@ III Evolution towards the minimum
$V_{\rm LVS}(\Phi)=V_{0}\left[(1-\epsilon\,\Phi^{3/2})\,e^{-3\sqrt{\frac{3}{2}}\,\Phi}+\delta\,e^{-2\sqrt{\frac{3}{2}}\,\Phi}\right],$ (9)
where the monomial term comes from integrating out the blow-up mode. This potential features a steep exponential behaviour on the left of a Minkowski (or de Sitter) minimum, followed by a relatively shallow maximum. This is at the core of the overshoot problem Brustein and Steinhardt (1993); Kaloper and Olive (1993); Barreiro et al. (1998); Brustein et al. (2004); Kaloper et al. (2005); Battefeld and Watson (2006); Conlon et al. (2008); Acharya et al. (2008); Conlon and Revello (2022); Mosny et al. (2025).
If the field reaches its minimum with too much kinetic energy, as would happen in a kination scenario, it would cross the maximum and inevitably roll towards infinity. This problem is therefore potentially present in all models with a kination epoch. However, as noted in Copeland et al. (1998), the kinating fixed point of the dynamical system of a scalar field with an exponential potential is unstable. In fact, if even a small percentage of energy density is in any background fluid, the system will move away from the kinating fixed point towards a scaling or tracker solution. Depending on the initial fraction of energy in the background fluid, the friction may decelerate the field enough to make it stop before it reaches the maximum. This is the typical solution to the overshoot problem, which works well for a field excursion of order $\Delta\phi\sim\mathcal{O}(10)\,M_{p}$ in the presence of a radiation background. The situation changes when we add a fluid of string loops. In fact, as noted in Sánchez González et al. (2025), this fluid redshifts as:
$\rho_{\rm loop}\sim a^{-3}\,\mu^{1/2}\,.$ (10)
@@
$\rho_{\rm loop}\sim a^{-9/2}\,$, (11)
i.e. the F-strings redshift faster than radiation, making up a rather peculiar fluid. However, the energy density in string loops depends on the behaviour of the scalar field through the field-dependence of the tension. Therefore, to understand how the loops influence the dynamics of the rolling modulus and vice-versa, a complete study of the dynamical system must be performed, including the contribution of the fluid of string loops. We shall perform this analysis combining the results of Sánchez González et al. (2025) with the requirement that the field does not overshoot its minimum at the end of its evolution. We shall study the system with and without radiation.
III.1 The dynamical system
The Friedman equation of a system formed by a scalar field $\phi$ with potential $V(\phi)$, loops of cosmic strings, and a fluid with energy density $\rho_{\rm f}$, can be written as (reintroducing $M_{p}$ to give a sense of the units):
$3H^{2}M_{p}^{2}=\rho_{\rm tot}=\frac{1}{2}\dot{\phi}^{2}+V(\phi)+\rho_{\rm loop}(\phi)+\rho_{\rm f}\,.$ (12)
@@ III.1 The dynamical system
$\dot{\rho}_{\rm loop}+3H\rho_{\rm loop}-\rho_{{\rm loop},\phi}\,\dot{\phi}=0\,.$ (15)
Differentiating (12) with respect to time, using the continuity equations and supposing the fluid is barotropic, $p_{\rm f}=\omega\,\rho_{\rm f}$, we get:
$2\dot{H}M_{p}^{2}=\dot{\phi}^{2}+\rho_{\rm loop}+(\omega+1)\rho_{\rm f}\,.$ (16)
@@
$X^{2}$ $\equiv\frac{\dot{\phi}^{2}}{2\rho_{\rm tot}}\equiv\Omega_{\rm k}\,$, (17)
$Y^{2}$ $\equiv\frac{V}{\rho_{\rm tot}}\equiv\Omega_{V}\,$, (18)
$Z^{2}$ $\equiv\frac{\rho_{\rm loop}}{\rho_{\rm tot}}\equiv\Omega_{\rm loop}\,$, (19)
$W^{2}$ $\equiv\frac{\rho_{\rm f}}{\rho_{\rm tot}}\equiv\Omega_{\rm f}\,$, (20)
@@
III.2 Case without radiation
@@
$Z^{2}=1-X^{2}-Y^{2}\,.$ (22)
Therefore, using the parametrisation (7) for the tension, we can write the autonomous system in the phase space variables $X$ and $Y$. To do that, we define:
$\lambda(\phi)\equiv-\sqrt{\frac{2}{3}}\frac{V,_{\phi}}{V}\,$, (23)
@@
$X^{\prime}$ $=\frac{3}{2}[X(X^{2}-Y^{2}-1)+\beta(1-X^{2})+(\lambda-\beta)Y^{2}]\,$, (24)
$Y^{\prime}$ $=\frac{3}{2}Y(X^{2}-Y^{2}-\lambda X+1)\,$, (25)
@@
The fixed points of this system for an exponential potential (i.e. $\lambda=\text{const}$) have been classified in Sánchez González et al. (2025). Here we report them in Tab. 1.
@@
$\mathcal{K}_{+}$ 1 0 0 $\qquad\qquad\qquad\qquad\quad\forall\lambda\,$ and $\forall\beta$
$\mathcal{K}_{-}$ -1 0 0 $\qquad\qquad\qquad\qquad\quad\forall\lambda\,$ and $\forall\beta$
$\mathcal{M}$ $\frac{\lambda}{2}$ $\sqrt{1-\frac{\lambda^{2}}{4}}$ 0 $\qquad\qquad\qquad\qquad\quad\;\forall\beta,\,\lambda\leq 2$
$\mathcal{L}$ $\beta$ 0 $\sqrt{1-\beta^{2}}$ $\qquad\qquad\qquad\qquad\quad\,\forall\lambda,\,\beta\leq 1$
$\mathcal{T}_{1}$ $\frac{1}{\lambda-\beta}$ $\frac{\sqrt{\beta^{2}+1-\lambda\beta}}{\lambda-\beta}$ $\frac{\sqrt{\lambda^{2}-2-\lambda\beta}}{\lambda-\beta}$ $\qquad\quad\beta\leq 1\;$ and $\frac{\beta}{2}+\sqrt{\frac{\beta^{2}}{4}+2}\leq\lambda\leq\beta+\frac{1}{\beta}$
Table 1: Fixed points (FP) of the autonomous system (24)-(25) with existence conditions.
@@
Figure 1: Stability map of the fixed points of the dynamical system (24)-(25) with exponential potential.
If we consider the full LVS potential (9), we lose the general validity of this map. However, for small values of the field $\Phi$, sufficiently far from the minimum $\Phi_{\rm min}$ where all the contributions compete, the potential can still be reliably approximated by an exponential:
$V\simeq V_{0}\,e^{-3\sqrt{\frac{3}{2}}\,\Phi}\,$, (26)
@@
$\mu\sim M_{s}^{2}=\frac{\sqrt{g_{s}}\,M_{p}^{2}}{4\pi\mathcal{V}}\,.$ (27)
Canonically normalising the volume as in (8), the tension becomes exponential in $\Phi$:
$\mu\sim M_{p}^{2}\,e^{-\sqrt{\frac{3}{2}}\,\frac{\Phi}{M_{p}}}\,.$ (28)
Comparing this expression with the general parametrisation of the tension given in (7), we find:
$\beta=1/2\,.$ (29)
If the field range for which the approximation (26) holds is sufficiently large, the system will evolve towards its final attractor which corresponds to the loop tracker $\mathcal{L}$, as can be seen from Fig. 1 with $\beta=1/2$ and $\lambda=3$. This is one of the main results of Sánchez González et al. (2025) which also found that in this scenario the loop energy density fraction grows significantly up to about $75\%$. However, as can be seen in Tab. 1, the kinetic energy of the field for $\beta=1/2$ is still $25\%$ of the total energy density, and so the rolling volume would generally overshoot the minimum. This is shown in Fig. 2. Here, the field $\Phi$ rolls down the full potential (9) in a background of loops of cosmic F-strings with various initial energy densities $Z_{0}^{2}\equiv\Omega_{\rm loop}^{(0)}$. In (9) we have chosen $V_{0}=1$, $\epsilon=0.013$ and $\delta\simeq 5.397\times 10^{-12}$ to obtain a Minkowski vacuum at $\Phi_{\rm min}=19\,M_{p}$ that would correspond, for $W_{0}\sim\mathcal{O}(10)$, to masses of the volume mode just above the bound of the cosmological moduli problem, $m_{\mathcal{V}}\sim W_{0}M_{p}/\mathcal{V}^{3/2}\gtrsim\mathcal{O}(30)$ TeV. It can be seen that the field overshoots the maximum of the potential, if released with vanishing velocity for a wide range of initial conditions.
Figure 2: Evolution of $\Phi$ towards the minimum of the potential (9) located at $\Phi_{\rm min}=19M_{p}$. For each initial value $\Phi_{0}$, 4 initial densities of loops of F-strings with $\beta=1/2$ have been tested: $Z_{0}^{2}=\{0.01,0.51,0.84,0.99\}$. All of them overshoot the minimum and the maximum (grey and black lines respectively).
Let us now investigate what happens when the fluid of cosmic string loops is made of effective strings. These are the result in the EFT of $p$-branes wrapping (p1)(p-1)-cycles in the internal dimensions, leaving one extended space dimension that looks like a string in spacetime. [^2] In Brunelli et al. (2025) these objects were called EFT strings. As shown in Revello and Villa (2025); Brunelli et al. (2025), the tension of such strings acquires an additional dependence on the volume of the wrapped cycle. In particular, if the wrapped (p1)(p-1)-cycle $\Sigma_{p-1}$ has volume $\text{Vol}(\Sigma_{p-1})$, then:
$\mu\sim M_{s}^{p+1}\,\text{Vol}(\Sigma_{p-1})\,.$ (30)
@@
    $\mu\sim M_{s}^{2}\,\mathcal{V}^{1/3}\sim M_{p}^{2}\,\mathcal{V}^{-2/3}\sim e^{-\sqrt{2/3}\,\Phi/M_{p}}\,.$ (31)
    Comparing with (7) we find:
    $\beta=1/3\,.$ (32)
@@
    $\mu\sim M_{s}^{2}\mathcal{V}^{2/3}\sim M_{p}^{2}\,\mathcal{V}^{-1/3}\sim e^{-1/\sqrt{6}\,\Phi/M_{p}}\,.$ (33)
    Comparing again with (7), we get:
    $\beta=1/6\,.$ (34)

### v2

@@ 1 Introduction
Early universe cosmology is a very promising arena to probe UV complete theories like string theory. This involves the study of trans-Planckian field excursions during epochs of large field inflation and kination, or the gravitational decay of long-lived moduli (see [14] for a recent review of different aspects of string cosmology).
@@
In this paper we focus on two features that typically arise at the end of string inflation (see for example brane-antibrane inflation [13, 25, 29, 7, 6, 16] or volume modulus inflation [22, 18]): a volume modulus that rolls over several Planck units towards its late time minimum, and a population of cosmic superstrings with a modulus-dependent tension that varies with time.
@@
A notorious issue of string models of the early universe is the overshoot problem [12], that is the tendency of the volume mode to run over the barrier that separates its minimum from the decompactification limit, if the modulus is released with a very high energy after the end of inflation. A known solution relies on the presence of a background fluid, like radiation or matter, which redshifts slower than the modulus kinetic energy and yields additional Hubble friction to the modulus dynamics [31, 4, 11, 32, 5, 28, 22, 1]. However, especially in models where the volume mode is rolling before reheating, motivating the microscopic origin of enough radiation to prevent overshooting can be non-trivial [23, 34].
@@
Recently, the powerful techniques of dynamical systems have been exploited to explore in detail the cosmological evolution of systems involving rolling scalars, cosmic superstrings and additional background fluids. Previous studies discovered that, during kination, the physical size of loops of fundamental strings (hereafter referred to as F-strings) [20] and effective strings arising from D3- and NS5-branes wrapped around fibration cycles [10] can grow faster than the scale factor and form a cosmic superstring network. Interestingly, this growth can occur for F-strings also for scaling fixed points in the presence of a background fluid with an appropriate equation of state parameter [10]. Moreover, ref. [36] took into account the energy density in cosmic string loops and revealed, among other things, the existence of a new attractor fixed point (a loop tracker) where the energy density of the universe is dominated by the cosmic superstrings.
@@
In this paper we extend the analysis of these cosmological systems considering the whole potential of the rolling volume modulus of type IIB LVS scenarios [3, 15], which includes both an early-time runaway and a late-time minimum with a barrier against decompactification. We find that, without enough radiation, a model with F-strings would face an overshoot problem. On the other hand, as soon as an initial population of effective strings arising from NS5-branes wrapped around 4-cycles is present, the overshoot problem is avoided, even in the absence of radiation. Moreover, the cosmic superstrings constitute around $97\%$ of the total energy density in the loop tracker fixed point that precedes the final evolution of the field towards the minimum of its potential where the energy density of cosmic superstrings can be as high as about $50\%$. Such large values of energy densities imply the potential generation of a detectable gravitational wave (GW) signal at high frequencies from the decay of cosmic superstrings in the early universe. The GW spectrum is expected to be peaked in the GHz regime and be diluted by the subsequent epoch of matter domination caused by the oscillations of the modulus around the minimum of its potential [21]. [^1] In string cosmology, there are several mechanisms for producing GWs, such as primordial tensor modes during inflation or secondary GWs sourced by large curvature or isocurvature perturbations; however, these typically lie at much lower frequencies.
2 Strings with time-dependent tension
@@
We start by giving a brief review of the derivation of the equations of motion for a string loop with time-dependent tension following [26, 20, 35]. The Nambu-Goto (NG) action for a string with a tension $\mu$ dependent on the spacetime coordinates $x^{\mu}$ is given by:
$S_{\rm NG}=-\int d^{2}\sigma\,\mu(x)\,\sqrt{-\gamma}\,.$ (2.1)
@@
$\gamma_{ab}=\frac{\partial x^{\mu}}{\partial\sigma^{a}}\frac{\partial x^{\nu}}{\partial\sigma^{b}}\,g_{\mu\nu}\,.$ (2.2)
Varying (2.1) with respect to $x^{\mu}$ yields the following equations of motion [26]:
$\frac{1}{\mu\sqrt{-\gamma}}\,\partial_{a}(\mu\sqrt{-\gamma}\gamma^{ab}x^{\mu}_{,b})+\Gamma^{\mu}_{\nu\rho}\gamma^{ab}x^{\nu}_{,a}x^{\rho}_{,b}-\frac{\partial^{\mu}\mu}{\mu}=0\,.$ (2.3)
@@
$\frac{\dot{\epsilon}}{\epsilon}$ $=H-a^{2}\dot{R}^{2}\left(2H+\frac{\dot{\mu}}{\mu}\right),$ (2.5)
$\ddot{R}+H\dot{R}\,+$ $\frac{R}{\epsilon^{2}}+\left(2H+\frac{\dot{\mu}}{\mu}\right)(1-a^{2}\dot{R}^{2})\dot{R}=0\,$, (2.6)
where $\epsilon\equiv a|R|/\sqrt{1-a^{2}\dot{R}^{2}}$ parametrizes the physical radius of the string at the oscillation turning points and is related to its (linear) energy density [33].
@@
$\mu(t)=\mu_{0}\,e^{-\sqrt{6}\beta\,\frac{\phi(t)}{M_{p}}}\ .$ (2.7)
We will first employ the techniques of [36] to study the cosmological evolution of a system including a population of cosmic superstrings and a modulus rolling towards the minimum of its potential, with and without radiation in generality. Then we will study what happens when the field stabilises at its minimum and the tension becomes oscillating, analysing the evolution of a single isolated string and whether it has any peculiar dynamics.
3 Evolution towards the minimum
All previous works on the evolution of a dilute gas of cosmic string loops assumed an exponential potential [20, 10, 36]. However this approximation holds only in a portion of a phenomenologically realistic potential. In particular, when the rolling field is the volume modulus of string compactifications, it must settle into a minimum at the end of its excursion to avoid a decompactification limit. The precise structure of the potential thus becomes relevant for assessing how the field reaches its minimum and whether it settles there without overshooting.
The potential we will focus on is the one of type IIB LVS models [3, 15]. This class of models features (at least) two moduli fields: one controlling the overall internal volume, and the other parametrising the size of a small blow-up cycle. At leading order, the LVS potential receives three types of contributions: $\mathcal{O}(\alpha^{\prime 3})$ corrections to the Kähler potential [8], non-perturbative effects in the superpotential [9], and a model-dependent uplift term [30, 19, 17]. After integrating out the blow-up mode, the potential depends only on the volume mode $\mathcal{V}$. The canonical normalisation of the volume is:
$\frac{\Phi}{M_{p}}=\sqrt{\frac{2}{3}}\ln\mathcal{V}\,.$ (3.1)
@@ 3 Evolution towards the minimum
$V_{\rm LVS}(\Phi)=V_{0}\left[(1-\epsilon\,\Phi^{3/2})\,e^{-3\sqrt{\frac{3}{2}}\,\Phi}+\delta\,e^{-2\sqrt{\frac{3}{2}}\,\Phi}\right],$ (3.2)
where the monomial term comes from integrating out the blow-up mode. This potential features a steep exponential behaviour on the left of a Minkowski (or de Sitter) minimum, followed by a relatively shallow maximum. This is at the core of the overshoot problem [12, 31, 4, 11, 32, 5, 22, 1, 23, 34].
If the field reaches its minimum with too much kinetic energy, as would happen in a kination scenario, it would cross the maximum and inevitably roll towards infinity. This problem is therefore potentially present in all models with a kination epoch. However, as noted in [24], the kinating fixed point of the dynamical system of a scalar field with an exponential potential is unstable. In fact, if even a small percentage of energy density is in any background fluid, the system will move away from the kinating fixed point towards a scaling or tracker solution. Depending on the initial fraction of energy in the background fluid, the friction may decelerate the field enough to make it stop before it reaches the maximum. This is the typical solution to the overshoot problem, which works well for a field excursion of order $\Delta\phi\sim\mathcal{O}(10)\,M_{p}$ in the presence of a radiation background. The situation changes when we add a fluid of string loops. In fact, as noted in [36], this fluid redshifts as:
$\rho_{\rm loop}\sim a^{-3}\,\mu^{1/2}\,.$ (3.3)
@@
$\rho_{\rm loop}\sim a^{-9/2}\,$, (3.4)
i.e. the F-strings redshift faster than radiation, making up a rather peculiar fluid. However, the energy density in string loops depends on the behaviour of the scalar field through the field-dependence of the tension. Therefore, to understand how the loops influence the dynamics of the rolling modulus and vice-versa, a complete study of the dynamical system must be performed, including the contribution of the fluid of string loops. We shall perform this analysis combining the results of [36] with the requirement that the field does not overshoot its minimum at the end of its evolution. We shall study the system with and without radiation.
3.1 The dynamical system
The Friedman equation of a system formed by a scalar field $\phi$ with potential $V(\phi)$, loops of cosmic strings, and a perfect fluid with energy density $\rho_{\rm f}$, can be written as (reintroducing $M_{p}$ to give a sense of the units):
$3H^{2}M_{p}^{2}=\rho_{\rm tot}=\frac{1}{2}\dot{\phi}^{2}+V(\phi)+\rho_{\rm loop}(\phi)+\rho_{\rm f}\,.$ (3.5)
@@ 3.1 The dynamical system
$\dot{\rho}_{\rm loop}+3H\rho_{\rm loop}-\rho_{{\rm loop},\phi}\,\dot{\phi}=0\,.$ (3.8)
Differentiating (3.5) with respect to time, using the continuity equations and supposing the fluid is barotropic, $p_{\rm f}=\omega\,\rho_{\rm f}$, we get:
$2\dot{H}M_{p}^{2}=\dot{\phi}^{2}+\rho_{\rm loop}+(\omega+1)\rho_{\rm f}\,.$ (3.9)
@@
$X^{2}$ $\equiv\frac{\dot{\phi}^{2}}{2\rho_{\rm tot}}\equiv\Omega_{\rm k}\,$, (3.10)
$Y^{2}$ $\equiv\frac{V}{\rho_{\rm tot}}\equiv\Omega_{V}\,$, (3.11)
$Z^{2}$ $\equiv\frac{\rho_{\rm loop}}{\rho_{\rm tot}}\equiv\Omega_{\rm loop}\,$, (3.12)
$W^{2}$ $\equiv\frac{\rho_{\rm f}}{\rho_{\rm tot}}\equiv\Omega_{\rm f}\,$, (3.13)
@@
3.2 Case without radiation
@@
$Z^{2}=1-X^{2}-Y^{2}\,.$ (3.15)
Therefore, using the parametrisation (2.7) for the tension, we can write the autonomous system in the phase space variables $X$ and $Y$. To do that, we define:
$\lambda(\phi)\equiv-\sqrt{\frac{2}{3}}\frac{V,_{\phi}}{V}\,$, (3.16)
@@
$X^{\prime}$ $=\frac{3}{2}[X(X^{2}-Y^{2}-1)+\beta(1-X^{2})+(\lambda-\beta)Y^{2}]\,$, (3.17)
$Y^{\prime}$ $=\frac{3}{2}Y(X^{2}-Y^{2}-\lambda X+1)\,$, (3.18)
@@
The fixed points of this system for an exponential potential (i.e. $\lambda=\text{const}$) have been classified in [36]. Here we report them in Tab. 1.
@@
$\mathcal{K}_{+}$ 1 0 0 $\qquad\qquad\qquad\quad\forall\lambda\,$ and $\forall\beta$
$\mathcal{K}_{-}$ -1 0 0 $\qquad\qquad\qquad\quad\forall\lambda\,$ and $\forall\beta$
$\mathcal{M}$ $\frac{\lambda}{2}$ $\sqrt{1-\frac{\lambda^{2}}{4}}$ 0 $\qquad\qquad\qquad\qquad\forall\beta,\,\lambda\leq 2$
$\mathcal{L}$ $\beta$ 0 $\sqrt{1-\beta^{2}}$ $\qquad\qquad\qquad\qquad\forall\lambda,\,\beta\leq 1$
$\mathcal{T}_{1}$ $\frac{1}{\lambda-\beta}$ $\frac{\sqrt{\beta^{2}+1-\lambda\beta}}{\lambda-\beta}$ $\frac{\sqrt{\lambda^{2}-2-\lambda\beta}}{\lambda-\beta}$ $\quad\beta\leq 1\;$ and $\frac{\beta}{2}+\sqrt{\frac{\beta^{2}}{4}+2}\leq\lambda\leq\beta+\frac{1}{\beta}$
Table 1: Fixed points (FP) of the autonomous system (3.17)-(3.18) with existence conditions.
@@
Figure 1: Stability map of the fixed points of the dynamical system (3.17)-(3.18) with exponential potential.
If we consider the full LVS potential (3.2), we lose the general validity of this map. However, for small values of the field $\Phi$, sufficiently far from the minimum $\Phi_{\rm min}$ where all the contributions compete, the potential can still be reliably approximated by an exponential:
$V\simeq V_{0}\,e^{-3\sqrt{\frac{3}{2}}\,\Phi}\,$, (3.19)
@@
$\mu\sim M_{s}^{2}=\frac{\sqrt{g_{s}}\,M_{p}^{2}}{4\pi\mathcal{V}}\,.$ (3.20)
Canonically normalising the volume as in (3.1), the tension becomes exponential in $\Phi$:
$\mu\sim M_{p}^{2}\,e^{-\sqrt{\frac{3}{2}}\,\frac{\Phi}{M_{p}}}\,.$ (3.21)
Comparing this expression with the general parametrisation of the tension given in (2.7), we find:
$\beta=1/2\,.$ (3.22)
If the field range for which the approximation (3.19) holds is sufficiently large, the system will evolve towards its final attractor which corresponds to the loop tracker $\mathcal{L}$, as can be seen from Fig. 1 with $\beta=1/2$ and $\lambda=3$. This is one of the main results of [36] which also found that in this scenario the loop energy density fraction grows significantly up to about $75\%$. However, as can be seen in Tab. 1, the kinetic energy of the field for $\beta=1/2$ is still $25\%$ of the total energy density, and so the rolling volume would generally overshoot the minimum. This is shown in Fig. 2. Here, the field $\Phi$ rolls down the full potential (3.2) in a background of loops of cosmic F-strings with various initial energy densities $Z_{0}^{2}\equiv\Omega_{\rm loop}^{(0)}$. In (3.2) we have chosen $V_{0}=1$, $\epsilon=0.013$ and $\delta\simeq 5.397\times 10^{-12}$ to obtain a Minkowski vacuum at $\Phi_{\rm min}=19\,M_{p}$ that would correspond, for $W_{0}\sim\mathcal{O}(10)$, to masses of the volume mode just above the bound of the cosmological moduli problem, $m_{\mathcal{V}}\sim W_{0}M_{p}/\mathcal{V}^{3/2}\gtrsim\mathcal{O}(30)$ TeV. It can be seen that the field overshoots the maximum of the potential, if released with vanishing velocity for a wide range of initial conditions.
Figure 2: Evolution of $\Phi$ towards the minimum of the potential (3.2) located at $\Phi_{\rm min}=19M_{p}$. For each initial value $\Phi_{0}$, 4 initial densities of loops of F-strings with $\beta=1/2$ have been tested: $Z_{0}^{2}=\{0.01,0.51,0.84,0.99\}$. All of them overshoot the minimum and the maximum (grey and black lines respectively).
Let us now investigate what happens when the fluid of cosmic string loops is made of effective strings. These are the result in the EFT of $p$-branes wrapping (p1)(p-1)-cycles in the internal dimensions, leaving one extended space dimension that looks like a string in spacetime. [^3] In [10] these objects were called EFT strings. As shown in [35, 10], the tension of such strings acquires an additional dependence on the volume of the wrapped cycle. In particular, if the wrapped (p1)(p-1)-cycle $\Sigma_{p-1}$ has volume $\text{Vol}(\Sigma_{p-1})$, then:
$\mu\sim M_{s}^{p+1}\,\text{Vol}(\Sigma_{p-1})\,.$ (3.23)
@@
    $\mu\sim M_{s}^{2}\,\mathcal{V}^{1/3}\sim M_{p}^{2}\,\mathcal{V}^{-2/3}\sim e^{-\sqrt{2/3}\,\Phi/M_{p}}\,.$ (3.24)
    Comparing with (2.7) we find:
    $\beta=1/3\,.$ (3.25)
@@
    $\mu\sim M_{s}^{2}\mathcal{V}^{2/3}\sim M_{p}^{2}\,\mathcal{V}^{-1/3}\sim e^{-1/\sqrt{6}\,\Phi/M_{p}}\,.$ (3.26)
    Comparing again with (2.7), we get:
    $\beta=1/6\,.$ (3.27)